solve by substitution method
5x+2y=-7
7x+3y=13
I am having an extremly hard time figuring this out PLEASE help

Answer by Alan3354(69443)

solve by substitution method
5x+2y=-7
7x+3y=13
----------
The usual approach is to solve for either x or y in 1 of the eqns.
----
5x+2y=-7
2y = -5x - 7
y = -5x/2 - 7/2
----
Sub for y in the other eqn
7x+3y=13
7x + 3(-5x/2 - 7/2) = 13
7x - 15x/2 - 21/2 = 13
-x/2 = 47/2
x = -47
-------
y = -5x/2 - 7/2
y = 235/2 - 7/2
y = 114
======================
A different method:
5x+2y=-7
7x+3y=13
----
15x + 6y = -21
14x + 6y = 26
Solve for 6y, then sub. You avoid fractions that way.
15x + 6y = -21
--> 6y = -15x - 21
--------------
21x + 6y = 39
14x + (-15x - 21) = 26
etc, same answer